# The Role of Video Generation in Enhancing Data-Limited Action Understanding

### v1

@@ 1 Introduction
Over the past decade, deep learning  lecun2015deep ; miech2019howto100m has brought remarkable progress benefiting from large-scale annotated data carreira2018short ; sultani2018real , especially in the computer vision community such as CLIP with 400M image-text pairs radford2021learning . However, obtaining high-quality datasets is often complex, time-consuming, and costly carreira2018short ; grauman2022ego4d , requiring a large amount of manual annotation. In particular in the field of video understanding, human annotations can be inaccurate (50% samples are not aligned in the HowTo100M dataset  miech2019howto100m ). A lack of a well-annotated dataset with sufficient samples will limit the learning of models and result in poor generalization ability, facing challenges in real-world applications such as abnormal action detection sultani2018real ; lv2023unbiased and long-tail action recognition perrett2023use ; grauman2022ego4d .
@@
In order to solve the data-limited problem in video understanding, previous methods often focus on designing data augmentation strategies  yun2020videomix ; kim2020learning ; li2022video and transfer knowledge  kim2025leveraging ; li2024large ; rasheed2023fine ; luo2023towards from other modalities, such as images, where data is easier to collect. However, we argue that these solutions are suboptimal as a) data augmentation does not create novel semantics for general knowledge learning trabuccoeffective , and b) knowledge learned from other modalities does not provide necessary information for video understanding, specifically, images do not include temporal relations between video frames.
@@
With the development of diffusion models  yang2024cogvideox ; liu2024sora ; esser2023structure , existing methods propose to synthesis datasets to improve model learning he2023synthetic ; tian2024stablerep ; luo2024generative ; feng2023diverse . There is a problem in the utilization of generated samples for training video understanding models, manifested by the deficient information within the samples, which leads to suboptimal training efficiency. As illustrated in Figure LABEL:fig:clust (a) and (b), our finding is that due to the complex and rich information of real videos, real samples have higher intra-class distances and lower inter-class distances than synthetic samples; while the videos generated by the diffusion transformer tend to include information with simple and similar content, which limits the learnable information contained in the synthetic samples . This also makes the classification difficulty of synthetic samples lower than that of real samples.
@@
To solve the problem of deficient information within generated samples, we propose an information enhancement strategy to enhance the effective information of synthetic samples, which refers to context information conducive to action understanding. This strategy injects various character and environment information that conforms to specific human actions into the generated video and alleviates the domain gap between synthetic and real samples. Due to the limited capabilities of the text2video model, some synthetic videos perform unsatisfactorily and contain less effective semantic information, as shown in Figure LABEL:fig:clust (c). To alleviate the impact of low-quality synthetic samples on model training, we propose uncertainty-based label smoothing. We calculate the uncertainty of the generated samples with CLIP to measure their quality and adjust the smoothness of label smoothing with uncertainty, which can prevent the model from overfitting low-quality samples and alleviating its impact.
@@ 2.1 Training with Synthetic Samples
The methods of training with synthetic samples can be broadly categorized into two main streams: generative-based and graphics engine-based methods. The graphics engine-based methods use real objects’ 3D models to arrange, reconstruct, or move them according to certain rules and then render them to generate training samples. It has found usage in a myriad of fields, including object detection gaidon2016virtual ; cabon2020virtual , optical flow estimation dosovitskiy2015flownet ; kim2022transferable , auto driving dosovitskiy2017carla , etc.
With the explosion development of AIGC, image generation models have become of sufficient quality to generate training samples. Earlier works on image understanding have concentrated on the generation of samples or data augmentation with GANs  baranchuk2021label ; li2022bigdatasetgan . But with the emergence of the diffusion model, many methods focus on generating samples using diffusion models given its realistic, high-quality and controllable generative ability. he2023synthetic found that diffusion model synthetic samples can significantly improve the classification results of zero-shot, few-shot and transfer learning, and designed strategies to filter the noise of synthetic samples and enhance their diversity. feng2023diverse proposes a novel test-time prompt tuning method which leverages diffusion models to generate augmented data. zhou2023training proposes diffusion inversion, which invert images to the diffusion model’s latent space, and generates novel samples by conditioning the generative model. feng2024instagen fine-tunes a diffusion model to enhance the capability of the object detection model. wanggenerated studies the reasons why synthetic samples sometimes harms contrastive learning performance from the perspective of data inflation and data augmentation, and proposes adaptive inflation to improve the contrastive learning. trabuccoeffective uses a diffusion model to edit and change the semantics of the image to address the lack of diversity in previous image augmentation methods.
Within the realm of video understanding, most works still rely on graphics engines or GANs guo2022learning ; kim2022transferable . Unlike these previous works, this paper focuses on leveraging text2video diffusion transformers to improve data-limited action understanding that includes action recognition and action detection tasks. To the best of our knowledge, no prior research has delved into this domain.
@@ 2.2 Text-to-Video Diffusion Models
In recent years, text-to-video generation has undergone significant advancements, particularly driven by the rapid development of diffusion models. Previous video generation models based on diffusion typically followed a UNet-based architecture. Video Diffusion Models (VDMs) ho2022video extend image synthesis diffusion models to video generation by training jointly on both image and video data. Gen-1 esser2023structure presents a structure and content-aware model that modifies videos guided by example images or texts. With the impressive capabilities demonstrated by Sora liu2024sora , which uses Transformer as the backbone of diffusion models, i.e. Diffusion Transformers (DiT), has gradually become mainstream. A series of DiT-based works have emerged, including CogVideoX yang2024cogvideox and Hunyuan Video kong2024hunyuanvideo . These works are capable of generating high-resolution videos with coherent actions, which makes it possible to apply the generated long-tail video data for downstream task training. In this paper, we utilize CogVideoX yang2024cogvideox to generate all training samples because it performs well in generating human-related videos and includes rich motion information.
@@ 3.1 Overview
We address the data-limited action understanding problem by bridging the data scarcity. To this end, we adopt the text2video model to generate the required data according to the target task. Our approach consists of two steps: 1. Generate a video dataset with a text2video model based on given labels; 2. Train the action understanding model with the generated video dataset, as shown in Figure LABEL:fig:overall (a). This paradigm is concise yet general and can produce infinite-scale annotated data without human intervention.
@@ 3.3 Information Enhancement Strategy
Given an action category name $c$, we construct the contextual information most relevant to this action. We mainly focus on information from two perspectives: the environment where the action occurs and the appearance characteristics of the person performing the action. We generate contextual information through GPT-4o, where we generate 4 different environment descriptions $Env=gpt(prompt_{env},c)$ and 16 different character descriptions $Char=gpt(prompt_{char},c)$ for each action. Finally, we use this contextual information to enhance the content of the generated video. The process of generating action description information is shown in Figure LABEL:fig:overall (b).
@@ 3.4 Uncertainty-Based Label Smoothing
Due to the limited capabilities of the text2video model, generated videos sometimes have suboptimal generation effects and contain less effective semantic information for training. Such samples may be detrimental to the training of the model, as shown in Figure LABEL:fig:clust.
To address this problem, we propose uncertainty-based label smoothing. Given a synthetic video sample $v^{\prime}$, we first use CLIP radford2021learning to calculate its similarity $S$ with all category names $C$, and then use the entropy of the similarity to measure its uncertainty.
$\begin{split}S=\{CLIP(v^{\prime},c)|c\in C\},\\ H(S)=\sum_{i=1}^{k}S_{i}log\,S_{i}\end{split}$ (3)
@@
whle $w$ represents the weighting of uncertainty-based label smoothing, $y=[q_{1},q_{2},...,q_{k-1},q_{k}]$ represents the original one-hot label. As low-quality generated samples often exhibit greater uncertainty, leading to an increased degree of label smoothing in such cases, as shown in Figure LABEL:fig:overall (c). In this way, we can prevent the model from overfitting to low-quality samples and mitigate their impact on model training.
@@ 4.1 Implementation Details
The text2video model we use is CogVideoX-2B yang2024cogvideox . For each dataset, we generate 128 videos for each category with 50 inference steps. We need to emphasize that the training data of CogVideoX-2B does not contain the annotated UCF-101, HMDB-51 or Kinetics-600 datasets we used in our experiments, so there is no data leakage. For information enhancement strategy, we use GPT-4o to generate multiple different environment descriptions at once, with the following prompt like “Given a phrase describing an action, generate four different scene appearance descriptions that are most suitable for performing the action.”. Similarly, the prompt for generating character descriptions is as follows: “Given a phrase describing an action, generate 16 different descriptions of the physical features of a person who is suitable for that action.”. In subsequent training, we adopt TC-CLIP kim2025leveraging in most experiments except for abnormal action detection. For abnormal action detection, we use the X-CLIP-B/32 ni2022expanding following lv2023unbiased . For tasks where real samples are not available such as zero-shot, we train the model with synthetic samples only. For tasks where real samples are available such as few-shot, long-tail, etc., we pre-train the model with synthetic samples and then fine-tune with the real samples. We set the $w$ to 0.3 in uncertainty-based label smoothing. After calculating the uncertainty of all samples in a dataset, normalization is performed on them to constrain the scale.
@@
Vanilla CLIP radford2021learning 40.8 ± 0.3 63.2 ± 0.2 59.8 ± 0.3 83.5 ± 0.2 54.6
ActionCLIP wang2021actionclip 49.1 ± 0.4 68.0 ± 0.9 56.1 ± 0.9 83.2 ± 0.2 57.7
X-CLIP ni2022expanding 44.6 ± 5.2 72.0 ± 2.3 65.2 ± 0.4 86.1 ± 0.8 60.6
Vita-CLIP wasim2023vita 48.6 ± 0.6 75.0 ± 0.6 67.4 ± 0.5 - 63.7
Open-VCLIP weng2023open 53.9 ± 1.2 83.4 ± 1.2 73.0 ± 0.8 93.2 ± 0.1 70.1
OTI  zhu2023orthogonal 54.2 ± 1.3 83.3 ± 0.3 66.9 ± 1.0 - 68.1
OST chen2024ost 55.9 ± 1.2 79.7 ± 1.1 75.1 ± 0.6 94.6 ± 0.2 70.2
FROSTER huangfroster 54.8 ± 1.3 84.8 ± 1.1 74.8 ± 0.9 - 71.5
IMP-MoE-L (VIT-L)  akbari2023alternating 59.1 91.5 76.8 - 75.8
OTI (VIT-L)  zhu2023orthogonal 59.3 ± 1.7 88.1 ± 1.0 70.6 ± 0.5 - 72.7
ViFi-CLIP rasheed2023fine 51.3 ± 0.6 76.8 ± 0.7 71.2 ± 1.0 92.2 ± 0.3 66.4
ViFi-CLIP rasheed2023fine +Ours 60.8 ± 0.4 81.9 ± 1.2 76.0 ± 1.0 94.3 ± 0.2 72.9
BIKE Wu_2023_CVPR 53.3 ± 1.1 79.6 ± 0.3 68.5 ± 1.3 90.9 ± 0.4 67.1
BIKE Wu_2023_CVPR +Ours 56.5 ± 0.7 87.7 ± 1.0 78.0 ± 1.1 94.7 ± 0.3 74.1
TC-CLIP kim2025leveraging 56.0 ± 0.3 85.4 ± 0.8 78.1 ± 1.0 95.7 ± 0.3 73.2
TC-CLIP kim2025leveraging +Ours 61.0 ± 1.0 86.5 ± 0.7 78.6 ± 0.8 96.1 ± 0.2 75.3
@@
Vanilla CLIP radford2021learning 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6
ActionCLIP wang2021actionclip 47.5 57.9 57.3 59.1 70.6 71.5 73.0 91.4
X-CLIP ni2022expanding 53.0 57.3 62.8 64.0 76.4 83.4 88.3 91.4
ViFi-CLIP rasheed2023fine 57.2 62.7 64.5 66.8 80.7 85.1 90.0 92.7
OST chen2024ost 59.1 62.9 64.9 68.2 82.5 87.5 91.7 93.9
TC-CLIP kim2025leveraging 58.6 63.3 65.5 68.8 86.8 90.1 92.0 94.3
TC-CLIP kim2025leveraging +Ours 63.0 65.6 68.9 71.4 88.1 91.3 92.9 94.0
@@
Vanilla CLIP radford2021learning 53.3 46.8 49.8 78.5 63.6 70.3
ActionCLIP wang2021actionclip 69.1 37.3 48.5 90.1 58.1 70.7
X-CLIP ni2022expanding 69.4 45.5 55.0 89.9 58.9 71.2
ViFi-CLIP rasheed2023fine 73.8 53.3 61.9 92.9 67.7 78.3
Open-VCLIP weng2023open 70.3 50.4 58.9 94.8 77.5 85.3
FROSTER huangfroster 74.1 58.0 65.1 95.3 80.0 87.0
TC-CLIP kim2025leveraging 73.3 59.1 65.5 95.4 81.6 88.0
TC-CLIP kim2025leveraging +Ours 77.7 64.0 70.2 95.5 85.2 90.1
@@
TC-CLIP kim2025leveraging 49.41 79.63 87.08 60.39 76.13 66.48 80.03 73.46
TC-CLIP kim2025leveraging +Ours 52.55 80.37 87.50 62.40 79.30 68.36 86.25 78.77
@@
Sohrab et al. sohrab2018subspace 58.50 -
BODS wang2019gods 68.26 -
GODS wang2019gods 70.46 -
Zhang et al. zhang2019temporal 78.66 -
Motion-Aware zhu2019motion 79.10 62.18
Wu et al. wu2020not 82.44 -

### v2

@@ 1 Introduction
Figure 1: The samples (a) and t-SNE visualizations (b) of the synthetic dataset and the real dataset. From left to right, they are: the HMDB-51 dataset, the synthetic HMDB-51 dataset with the basic strategy and the synthetic HMDB-51 dataset with our proposed information enhancement strategy. (c) Unsatisfactory synthetic videos.
@@
Over the past decade, deep learning  LeCun et al. (2015); Miech et al. (2019) has brought remarkable progress benefiting from large-scale annotated data Carreira et al. (2018); Sultani et al. (2018), especially in the computer vision community such as CLIP with 400M image-text pairs Radford et al. (2021). However, obtaining high-quality datasets is often complex, time-consuming, and costly Carreira et al. (2018); Grauman et al. (2022), requiring a large amount of manual annotation. In particular in the field of video understanding, human annotations can be inaccurate (50% samples are not aligned in the HowTo100M dataset  Miech et al. (2019)). A lack of a well-annotated dataset with sufficient samples will limit the learning of models and result in poor generalization ability, facing challenges in real-world applications such as abnormal action detection Sultani et al. (2018); Lv et al. (2023) and long-tail action recognition Perrett et al. (2023); Grauman et al. (2022).
@@
In order to solve the data-limited problem in video understanding, previous methods often focus on designing data augmentation strategies  Yun et al. (2020); Kim et al. (2020); Li et al. (2022b) and transfer knowledge  Kim et al. (2025); Li et al. (2024); Rasheed et al. (2023); Luo et al. (2023) from other modalities, such as images, where data is easier to collect. However, we argue that these solutions are suboptimal as a) data augmentation does not create novel semantics for general knowledge learning Trabucco et al. (2024), and b) knowledge learned from other modalities does not provide necessary information for video understanding, specifically, images do not include temporal relations between video frames.
@@
With the development of diffusion models  Yang et al. (2024); Liu et al. (2024); Esser et al. (2023), existing methods propose to synthesis datasets to improve model learning He et al. (2023); Tian et al. (2024); Luo et al. (2024); Feng et al. (2023). There is a problem in the utilization of generated samples for training video understanding models, manifested by the deficient information within the samples, which leads to suboptimal training efficiency. As illustrated in Figure 1 (a) and (b), our finding is that due to the complex and rich information of real videos, real samples have higher intra-class distances and lower inter-class distances than synthetic samples; while the videos generated by the diffusion transformer tend to include information with simple and similar content, which limits the learnable information contained in the synthetic samples . This also makes the classification difficulty of synthetic samples lower than that of real samples.
@@
To solve the problem of deficient information within generated samples, we propose an information enhancement strategy to enhance the effective information of synthetic samples, which refers to context information conducive to action understanding. This strategy injects various character and environment information that conforms to specific human actions into the generated video and alleviates the domain gap between synthetic and real samples. Due to the limited capabilities of the text2video model, some synthetic videos perform unsatisfactorily and contain less effective semantic information, as shown in Figure 1 (c). To alleviate the impact of low-quality synthetic samples on model training, we propose uncertainty-based label smoothing. We calculate the uncertainty of the generated samples with CLIP to measure their quality and adjust the smoothness of label smoothing with uncertainty, which can prevent the model from overfitting low-quality samples and alleviating its impact.
@@ 2.1 Training with Synthetic Samples
The methods of training with synthetic samples can be broadly categorized into two main streams: generative-based and graphics engine-based methods. The graphics engine-based methods use real objects’ 3D models to arrange, reconstruct, or move them according to certain rules and then render them to generate training samples. It has found usage in a myriad of fields, including object detection Gaidon et al. (2016); Cabon et al. (2020), optical flow estimation Dosovitskiy et al. (2015); Kim et al. (2022), auto driving Dosovitskiy et al. (2017), etc.
With the explosion development of AIGC, image generation models have become of sufficient quality to generate training samples. Earlier works on image understanding have concentrated on the generation of samples or data augmentation with GANs  Baranchuk et al. (2021); Li et al. (2022a). But with the emergence of the diffusion model, many methods focus on generating samples using diffusion models given its realistic, high-quality and controllable generative ability. He et al. (2023) found that diffusion model synthetic samples can significantly improve the classification results of zero-shot, few-shot and transfer learning, and designed strategies to filter the noise of synthetic samples and enhance their diversity. Feng et al. (2023) proposes a novel test-time prompt tuning method which leverages diffusion models to generate augmented data. Zhou et al. (2023) proposes diffusion inversion, which invert images to the diffusion model’s latent space, and generates novel samples by conditioning the generative model. Feng et al. (2024) fine-tunes a diffusion model to enhance the capability of the object detection model. Wang et al. (2024) studies the reasons why synthetic samples sometimes harms contrastive learning performance from the perspective of data inflation and data augmentation, and proposes adaptive inflation to improve the contrastive learning. Trabucco et al. (2024) uses a diffusion model to edit and change the semantics of the image to address the lack of diversity in previous image augmentation methods.
Within the realm of video understanding, most works still rely on graphics engines or GANs Guo et al. (2022); Kim et al. (2022). Unlike these previous works, this paper focuses on leveraging text2video diffusion transformers to improve data-limited action understanding that includes action recognition and action detection tasks. To the best of our knowledge, no prior research has delved into this domain.
@@ 2.2 Text-to-Video Diffusion Models
In recent years, text-to-video generation has undergone significant advancements, particularly driven by the rapid development of diffusion models. Previous video generation models based on diffusion typically followed a UNet-based architecture. Video Diffusion Models (VDMs) Ho et al. (2022) extend image synthesis diffusion models to video generation by training jointly on both image and video data. Gen-1 Esser et al. (2023) presents a structure and content-aware model that modifies videos guided by example images or texts. With the impressive capabilities demonstrated by Sora Liu et al. (2024), which uses Transformer as the backbone of diffusion models, i.e. Diffusion Transformers (DiT), has gradually become mainstream. A series of DiT-based works have emerged, including CogVideoX Yang et al. (2024) and Hunyuan Video Kong et al. (2024). These works are capable of generating high-resolution videos with coherent actions, which makes it possible to apply the generated long-tail video data for downstream task training. In this paper, we utilize CogVideoX Yang et al. (2024) to generate all training samples because it performs well in generating human-related videos and includes rich motion information.
@@ 3 Method
Figure 2: (a): The overall structure of our proposed method. We designed two strategies for generating sample training, information enhancement strategy (left) and uncertainty-based label smoothing strategy (right). (b): The process of generating action description information through proposed information enhancement strategy. (c): Uncertainty-based label smoothing uses a higher smoothness for low-quality generated samples with higher uncertainty.
@@ 3.1 Overview
We address the data-limited action understanding problem by bridging the data scarcity. To this end, we adopt the text2video model to generate the required data according to the target task. Our approach consists of two steps: 1. Generate a video dataset with a text2video model based on given labels; 2. Train the action understanding model with the generated video dataset, as shown in Figure 2 (a). This paradigm is concise yet general and can produce infinite-scale annotated data without human intervention.
@@ 3.3 Information Enhancement Strategy
Given an action category name $c$, we construct the contextual information most relevant to this action. We mainly focus on information from two perspectives: the environment where the action occurs and the appearance characteristics of the person performing the action. We generate contextual information through GPT-4o, where we generate 4 different environment descriptions $Env=gpt(prompt_{env},c)$ and 16 different character descriptions $Char=gpt(prompt_{char},c)$ for each action. Finally, we use this contextual information to enhance the content of the generated video. The process of generating action description information is shown in Figure 2 (b).
@@ 3.4 Uncertainty-Based Label Smoothing
Due to the limited capabilities of the text2video model, generated videos sometimes have suboptimal generation effects and contain less effective semantic information for training. Such samples may be detrimental to the training of the model, as shown in Figure 1.
To address this problem, we propose uncertainty-based label smoothing. Given a synthetic video sample $v^{\prime}$, we first use CLIP Radford et al. (2021) to calculate its similarity $S$ with all category names $C$, and then use the entropy of the similarity to measure its uncertainty.
$\begin{split}S=\{CLIP(v^{\prime},c)|c\in C\},\\ H(S)=-\sum_{i=1}^{k}S_{i}log\,S_{i}\end{split}$ (3)
@@
whle $w$ represents the weighting of uncertainty-based label smoothing, $y=[q_{1},q_{2},...,q_{k-1},q_{k}]$ represents the original one-hot label. As low-quality generated samples often exhibit greater uncertainty, leading to an increased degree of label smoothing in such cases, as shown in Figure 2 (c). In this way, we can prevent the model from overfitting to low-quality samples and mitigate their impact on model training.
@@ 4.1 Implementation Details
The text2video model we use is CogVideoX-2B Yang et al. (2024). For each dataset, we generate 128 videos for each category with 50 inference steps. We need to emphasize that the training data of CogVideoX-2B does not contain the annotated UCF-101, HMDB-51 or Kinetics-600 datasets we used in our experiments, so there is no data leakage. For information enhancement strategy, we use GPT-4o to generate multiple different environment descriptions at once, with the following prompt like “Given a phrase describing an action, generate four different scene appearance descriptions that are most suitable for performing the action.”. Similarly, the prompt for generating character descriptions is as follows: “Given a phrase describing an action, generate 16 different descriptions of the physical features of a person who is suitable for that action.”. In subsequent training, we adopt TC-CLIP Kim et al. (2025) in most experiments except for abnormal action detection. For abnormal action detection, we use the X-CLIP-B/32 Ni et al. (2022) following Lv et al. (2023). For tasks where real samples are not available such as zero-shot, we train the model with synthetic samples only. For tasks where real samples are available such as few-shot, long-tail, etc., we pre-train the model with synthetic samples and then fine-tune with the real samples. We set the $w$ to 0.3 in uncertainty-based label smoothing. After calculating the uncertainty of all samples in a dataset, normalization is performed on them to constrain the scale.
@@
Vanilla CLIP Radford et al. (2021) 40.8 ± 0.3 63.2 ± 0.2 59.8 ± 0.3 83.5 ± 0.2 54.6
ActionCLIP Wang et al. (2021) 49.1 ± 0.4 68.0 ± 0.9 56.1 ± 0.9 83.2 ± 0.2 57.7
X-CLIP Ni et al. (2022) 44.6 ± 5.2 72.0 ± 2.3 65.2 ± 0.4 86.1 ± 0.8 60.6
Vita-CLIP Wasim et al. (2023) 48.6 ± 0.6 75.0 ± 0.6 67.4 ± 0.5 - 63.7
Open-VCLIP Weng et al. (2023) 53.9 ± 1.2 83.4 ± 1.2 73.0 ± 0.8 93.2 ± 0.1 70.1
OTI  Zhu et al. (2023) 54.2 ± 1.3 83.3 ± 0.3 66.9 ± 1.0 - 68.1
OST Chen et al. (2024) 55.9 ± 1.2 79.7 ± 1.1 75.1 ± 0.6 94.6 ± 0.2 70.2
FROSTER Huang et al. (2024) 54.8 ± 1.3 84.8 ± 1.1 74.8 ± 0.9 - 71.5
IMP-MoE-L (VIT-L)  Akbari et al. (2023) 59.1 91.5 76.8 - 75.8
OTI (VIT-L)  Zhu et al. (2023) 59.3 ± 1.7 88.1 ± 1.0 70.6 ± 0.5 - 72.7
ViFi-CLIP Rasheed et al. (2023) 51.3 ± 0.6 76.8 ± 0.7 71.2 ± 1.0 92.2 ± 0.3 66.4
ViFi-CLIP Rasheed et al. (2023)+Ours 60.8 ± 0.4 81.9 ± 1.2 76.0 ± 1.0 94.3 ± 0.2 72.9
BIKE Wu et al. (2023) 53.3 ± 1.1 79.6 ± 0.3 68.5 ± 1.3 90.9 ± 0.4 67.1
BIKE Wu et al. (2023)+Ours 56.5 ± 0.7 87.7 ± 1.0 78.0 ± 1.1 94.7 ± 0.3 74.1
TC-CLIP Kim et al. (2025) 56.0 ± 0.3 85.4 ± 0.8 78.1 ± 1.0 95.7 ± 0.3 73.2
TC-CLIP Kim et al. (2025)+Ours 61.0 ± 1.0 86.5 ± 0.7 78.6 ± 0.8 96.1 ± 0.2 75.3
@@
Vanilla CLIP Radford et al. (2021) 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6
ActionCLIP Wang et al. (2021) 47.5 57.9 57.3 59.1 70.6 71.5 73.0 91.4
X-CLIP Ni et al. (2022) 53.0 57.3 62.8 64.0 76.4 83.4 88.3 91.4
ViFi-CLIP Rasheed et al. (2023) 57.2 62.7 64.5 66.8 80.7 85.1 90.0 92.7
OST Chen et al. (2024) 59.1 62.9 64.9 68.2 82.5 87.5 91.7 93.9
TC-CLIP Kim et al. (2025) 58.6 63.3 65.5 68.8 86.8 90.1 92.0 94.3
TC-CLIP Kim et al. (2025)+Ours 63.0 65.6 68.9 71.4 88.1 91.3 92.9 94.0
@@
Vanilla CLIP Radford et al. (2021) 53.3 46.8 49.8 78.5 63.6 70.3
ActionCLIP Wang et al. (2021) 69.1 37.3 48.5 90.1 58.1 70.7
X-CLIP Ni et al. (2022) 69.4 45.5 55.0 89.9 58.9 71.2
ViFi-CLIP Rasheed et al. (2023) 73.8 53.3 61.9 92.9 67.7 78.3
Open-VCLIP Weng et al. (2023) 70.3 50.4 58.9 94.8 77.5 85.3
FROSTER Huang et al. (2024) 74.1 58.0 65.1 95.3 80.0 87.0
TC-CLIP Kim et al. (2025) 73.3 59.1 65.5 95.4 81.6 88.0
TC-CLIP Kim et al. (2025)+Ours 77.7 64.0 70.2 95.5 85.2 90.1
@@
TC-CLIP Kim et al. (2025) 49.41 79.63 87.08 60.39 76.13 66.48 80.03 73.46
TC-CLIP Kim et al. (2025)+Ours 52.55 80.37 87.50 62.40 79.30 68.36 86.25 78.77
@@
Sohrab et al. Sohrab et al. (2018) 58.50 -
BODS Wang and Cherian (2019) 68.26 -
GODS Wang and Cherian (2019) 70.46 -
Zhang et al. Zhang et al. (2019) 78.66 -
Motion-Aware Zhu and Newsam (2019) 79.10 62.18
Wu et al. Wu et al. (2020) 82.44 -